OpenStudy (anonymous):

A system of equations is shown below. a. Find the solution to the system of equations using either substitution or the linear combination method. Show all your work. i will put the numbers once i post this.

OpenStudy (anonymous):



OpenStudy (mrnood):

to use substitution you need to get one of the variables 'isolated' can oyu rewrite the second equation so it looks like y = ????

OpenStudy (anonymous):

I'll try. Give me a minute.

OpenStudy (mrnood):

hint: subtract 2x from each side....

OpenStudy (anonymous):

So it would be 2x + y = -34 -2x. -2x Y= 32

OpenStudy (mrnood):

no you have simply subtracted 2 from the right side you need to subtract 2x (x will still be in the equation but you have y on the left side alone...)

OpenStudy (anonymous):

So it would be x+y=-32

OpenStudy (mrnood):

no you started with -32 on the right side so what do you get on the right side if you subtract 2x from -32?

OpenStudy (anonymous):

-32x? Sorry im not good at algebra.

OpenStudy (mrnood):

I'm sorry - I think you need to be able to do this type of equation BEFORE you try to do systems of equations any term with x in it has to remain independent of the constant term so on the right side if you subtract 2x from -32 yo get -32 - 2x I'm afraid you probably are not ready to do the rest of this question

OpenStudy (anonymous):

Well this is on my home work and i need to try and get the answer. And after i finish school i wont need need this kind of math anyway so i dont bother to remember it.

OpenStudy (tkhunny):

Hogwash. Part of the learning of mathematics is thinking logically. You WILL need that ability. Part of the learning of mathematics is self confidence. You WILL need that. Part of the learning of mathematics is understanding the world of liars that waits to make you a victim because they know you are ignorant. You WILL need that defense. I could go on.

OpenStudy (anonymous):

I understand that but i dont need algebra for anything in my future. The life skills it teaches yes but not the concept.

OpenStudy (tkhunny):

You're missing the point. There are a thousand more examples. You WILL need algebra.